Circular No. 7346

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/2000 A1 (MONTANI) J. Montani, Lunar and Planetary Laboratory, reports his discovery of

a faint comet on CCD images taken with the 0.90-m Spacewatch telescope at Kitt Peak (discovery observation below). The comet shows a coma with diameter 5′′–6′′, slightly elongated in p.a. 245^{o}–250^{o}. An R CCD image taken by S. Kern with the 2.3-m Steward telescope on Jan. 13 shows the comet to be clearly extended toward the southwest, and she derived mag 18.1. W. Shook found the object to be nonstellar with a 2′′.6 tail toward the southwest on an image taken with the 3.5-m WIYN telescope on Jan. 13. Full astrometry and parabolic orbital elements appear on MPEC 2000-A46.

2000 UT  α_{2000}  δ_{2000}  m_{1}
Jan. 12.32564  8^{h}02^{m}08.^{s}15  +23^{o}15′21′′.2  18.9

SUPERNOVAE 1999gt AND 1999gu E. Cappellaro, A. Pastorello, L. Rizzi, M. Salvo, and M. Turatto,

Padua; J. Danziger and P. Mazzali, Trieste; F. Patat, European Southern Observatory (ESO); and S. Benetti, Telescopio Nazionale Galileo, La Palma, report the discovery of two supernovae on a CCD frame taken on 1999 Dec. 29.17 UT with the ESO/MPI 2.2-m telescope (+ Wide Field Imager) and centered on the AXAF deep field. SN 1999gt (V = 22.1) is located at α = 3^{h}32^{m}10^{s}.28, δ = −28^{o}06′16′′.0 (equinox 2000.0), which is 1′′.4 west of the host galaxy’s nucleus. SN 1999gu (V ∼ 22.2) is located at α = 3^{h}33^{m}00^{s}.15, δ = −27^{o}51′40′′.1, which is 0′′.4 east and 2′′.4 south of the host galaxy’s nucleus. The two supernovae were not present on a frame taken on Dec. 13.20 (limiting mag V = 23.5). Spectra (range 600–1000 nm, resolution 2 nm) of the new objects were obtained with the ESO 3.6-m telescope (+ EFOSC2) on 2000 Jan. 11.10 by P. Leisy and E. Pompei (ESO). The spectrum of the parent galaxy of SN 1999gt shows Hα, Hβ, and [O III] emission lines, providing a redshift of 0.274. Given this redshift, the spectrum of SN 1999gt resembles that of a type-Ia supernova, 1–2 weeks past maximum, though the Si II 615-nm absorption is not prominent. The spectrum of SN 1999gu shows a broad (FWHM = 9000 km/s) emission at 750 nm and an associated blue-shifted absorption. This feature is identified with the Hα, P-Cyg profile emission of a type-II supernova shortly after maximum. The expansion velocity of the ejecta is 10 000 km/s. On top of this line, a spatially extended, narrow Hα emission from the parent galaxy gives a redshift of 0.147.

2000 January 14  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green